# GeoTexDensifier: Geometry-Texture-Aware Densification for High-Quality Photorealistic 3D Gaussian Splatting

### v1

@@ Abstract
Index Terms:
3D Gaussian Splatting, adaptive density control, texture-aware densification, geometry-aware splitting.
@@ I Introduction
To better tackle both densification and positioning problems mentioned above, this paper presents a novel 3D Gaussian splatting framework for high-quality photorealistic novel view rendering, which we named GeoTexDensifier. Our framework innovatively proposes a densification strategy that enforces the reconstructed Gaussians to more accurately comply with the actual geometric structure of the scene. Moreover, our densified Gaussian model contains a sufficient number of splats to support fully textured regions, while maintaining sparse Gaussian splat distribution in weakly textured areas. In these ways, our novel 3DGS framework is able to produce high-quality 3D Gaussian models, as shown in the comparative example of “Church” in Fig. 1. The original 3DGS [1] has “over-reconstruction” problem that causes missing details in textured regions such as the carpet, while state-of-the-art (SOTA) works like GeoGaussian [15] improves the spatial structure of splat distribution and Pixel-GS [17] further densifies the splats in some textured regions like the carpet, but both of them still lack sufficient splat sampling to support texture details. Mini-Splatting [14] has sufficiently sampled splats but introduces over-densification situation in textureless regions, which might produce extra noisy Gaussians due to underconstrained visual ambiguity. In comparison, our GexTexDensifier delivers more accurate spatial distribution of Gaussian splat with less noise and more photorealistic novel view rendering effects according to the quantitative evaluation of the rendering results in the ground truth (GT) testing views on metrics of Structural SIMilarity (SSIM), Peak Signal to Noise Ratio (PSNR) in dB, and Learned Perceptual Image Patch Similarity (LPIPS). Our GeoTexDensifier framework carries out a geometry-aware splitting strategy to guide the splitting position sampling to fit the actual geometric surfaces of the scene with the help of normal prior, and filter out noisy splats with improper initial positions according to our Validation of Depth Ratio Change checking on relative depths provided by monocular depth prior. Meanwhile, a texture-aware densification strategy is adopted as auxiliary supplement for splitting of additional large Gaussians to produce a denser distribution of splats in fully textured areas, while keeping sparse distribution in textureless regions to maintain a high-quality Gaussian point cloud. The texture-aware densification and geometry-aware splitting strategies are fully combined in the iterative optimization to get the final high-quality photorealistic 3DGS model. Experiments on Mip-NeRF 360 [6], Tanks and Temples [16] datasets and self-captured scenes verify the effectiveness and robustness of our GeoTexDensifier pipeline.
@@ II-C Gaussian Optimization with Geometry Prior
Original 3DGS expects Gaussian splats expected to grow along the real scene surfaces, which is however not always conformed to. Consequently, some approaches have attempted to optimize Gaussians by leveraging geometry priors such as depth and normal maps. DN-Splatter [33] adopts a method similar to MonoSDF [34], by using mono-depths and mono-normals to constrain the rendering loss. DNGaussian [35] introduces Hard and Soft Depth Regularization to optimize the positions and opacities of Gaussians, and employs Global-Local Depth Normalization to mitigate the sensitivity of scale-invariant loss to small depth errors. PGSR [36] optimizes Gaussians by rendering unbiased depths combined with single-view and multi-view geometric consistency losses. Besides applying depth and normal constraints, both GaussianRoom [37] and GSDF [38] utilize a learnable neural SDF field to guide the growth of splats, so as to simultaneously optimize both SDF and Gaussians. These geometry prior based methods rely on the geometric accuracy of the prior guidance, and are easily influenced by the incorrect depths or normals, especially when the truth depth scale is unavailable, while our approach only uses relative depth ratios for verifying newly generated splats, so as to relax the requirement on depth or scale accuracy.
@@ III System Overview
Suppose each scene has a set of multi-view RGB images captured with digital cameras as input, denoted by =$\mathcal{I}=\{I_{t}|t=1,\dots,M\}$, where $M$ is the number of input images. Our GeoTexDensifier system is applied to the multi-view images as training views to robustly reconstruct an high-quality 3DGS model of the captured scene, which we defined as $\mathcal{G}$. Fig. 2 outlines the proposed high-quality 3DGS framework. In the initialization stage, COLMAP [39] is carried out first to estimate camera poses of all the input views denoted as $\{\mathbf{M}_{t}|t=1,\dots,M\}$, with sparse Structure-from-Motion (SfM) feature map points used to initialize a set of Gaussian splats. Meanwhile, a depth map is estimated for each input RGB image by ZoeDepth [40] with its normal map computed from depths, denoted as $\mathcal{D}=\{D_{t}|t=1,\dots,M\}$ and $\mathcal{N}=\{N_{t}|t=1,\dots,M\}$ respectively. After initialization finishes, the adaptive optimization stage follows to control the splat distribution by densification and pruning, while iteratively refining the Gaussian parameters to enforce their rasterized images to be consistent with the texture details observed in the training views. We use a texture-aware densification strategy to provide sufficient Gaussian splats in fully textured areas while maintaining the sparsity of Gaussians in regions with weak textures to ensure more accurate spatial distribution of Gaussian points. Additionally, the estimated depths are used as priors to guide the splitting of Gaussian splats to better conform to the actural tangential directions of the geometric surfaces in weakly textured regions and avoid the generation of isolated or scattered Gaussians caused by the ambiguity due to lack of visual features. The texture-aware densification approach together with the geometry-guided splitting strategy constitute our geometry-texture-aware densification framework for reconstructing high-quality photorealistic 3DGS model, which will be described in detail in the following sections.
@@ IV Preliminary
$&{\mathrm{C}}(\bar{{\mathbf{x}}})=\sum_{k=1}^{K}\mathrm{c}_{k}w_{k}(\bar{{\mathbf{x}}})\prod_{j=1}^{k-1}\big{(}1-w_{j}(\bar{{\mathbf{x}}}))\\ &w_{k}(\bar{{\mathbf{x}}})=\alpha_{k}\bar{G}_{k}(\bar{{\mathbf{x}}}),\end{array}$ (1)
where ${\mathrm{C}}(\bar{{\mathbf{x}}})$ represents the color rendered at image pixel $\bar{{\mathbf{x}}}$, and $w_{k}(\bar{{\mathbf{x}}})$ is the rendering weight for $\alpha$-blending. $\bar{G}_{k}\sim\mathcal{N}(\bar{\mu}_{k},\bar{\Sigma}_{k})$ is the projected 2D Gaussian distribution of $G_{k}$ through a local affine approximation of perspective transformation $\mathbf{W}$, defined as $\bar{\mu}_{k}={\mathbf{W}}\mu_{k}$ and $\bar{\Sigma}_{k}=[{\mathbf{J}}{\mathbf{W}}\Sigma_{k}{\mathbf{W}}^{\top}{\mathbf{J}}^{\top}]_{1:2,1:2}$, with $\mathbf{J}$ the Jacobian form of $\mathbf{W}$ and $[\cdot]_{1:2,1:2}$ taking the first two rows and columns as sub-matrix. Finally, by leveraging the differentiable rasterizer and comparing the rendered images to the training views, all attributes of the 3D Gaussians can be learned and optimized by minimizing the ${\mathcal{L}}_{1}$ loss combined with a D-SSIM loss.
@@
$&{\mathcal{G}}_{b}^{t}=\{G_{i}|S_{i}^{t}>{\mathcal{T}}\}\\ &S_{i}^{t}=\sum_{\bar{{\mathbf{x}}}=(1,1)}^{(W,H)}\delta(\mathbf{I}_{i}(\bar{{\mathbf{x}}})=\mathbf{I}_{max}(\bar{{\mathbf{x}}})).\end{array}$ (2)
@@ V Texture-Aware Densification
An ideal densification strategy is to generate splat distribution with a higher density in strongly textured areas, while still leaving Gaussian splats relatively sparse in more weakly textured regions such as surfaces with pure colors. The blur split strategy introduced by Mini-Splatting [14] is adopted in our ADC process, which is although helpful for densification, but might cause the following side effects which do not actually conform to our ideal purpose: firstly, weakly textured regions often have Gaussians with large contribution areas, leading to over-splitting of splats to consequently generate too many inadequately optimized floaters which will influence the rendering results, and the second is that an excessive number of Gaussians in uniform splat distribution over the whole scene will cost too heavy memory consumption and large output file storage.
@@
which effectively reflects the contribution to whether a Gaussian point should be split by a hyperbolic tangent function normalized to $[0,1]$, which is activated by the image gradient whose gray values are normalized from $[0,255]$ to $[0,1]$, as shown in Fig. 3(a). $\alpha_{s}$ and $\beta_{s}$ are the linear regression coefficiencies which we set to $20$ and $0.16$ respectively for our experiments. Fig. 3(c) shows a computed weight map by Eq. 3. With the newly defined weight, we can redefine the maximal contribution area of each Gaussian $G_{i}$ to $I_{t}$ as $\hat{S}_{i}^{t}$, by modifying Eq. 2 to:
$&{\mathcal{G}}_{b}^{t}=\{G_{i}|\hat{S}_{i}^{t}>\hat{{\mathcal{T}}}\}\\ &\hat{S}_{i}^{t}=\sum_{\bar{{\mathbf{x}}}=(1,1)}^{(W,H)}s_{t}(\bar{{\mathbf{x}}})\delta(\mathbf{I}_{i}(\bar{{\mathbf{x}}})=\mathbf{I}_{max}(\bar{{\mathbf{x}}})),\end{array}$ (4)
@@
This texture-aware densification serves as an auxiliary strategy that supplements the original splitting strategy for a better ADC.
@@ VI-B VDRC for Parent Splats
Before each new time of splitting, a certain number of iterations might shift some parent splats to improper locations. For example, $G_{a}$ might move from the door towards the wall, becoming an unreliable Gaussian like $G_{b}$, and will continue to generate a series of improper splats which will all meet the condition of VDRC. To better solve this issue, we apply VDRC to the parent Gaussians before each time we split them. Since the initial Gaussians are inherited from the SfM map points, we filter out the map points whose track lengths are less than $3$ or reprojection errors exceed $1$ pixel, to ensure the correctness of the initial splat locations. We also choose the image view with the smallest reprojection error denoted by $\widehat{V}(a)$ as the initial reference view of each Gaussian $G_{a}$, and add two additional attributes to record $\widehat{V}(a)$ and its initial depth ratio ^$\widehat{{\mathcal{R}}}_{\widehat{V}(a)}(G_{a})$. Before each time of splitting, each optimized parent Gaussian splat $G_{a}$ involved will be projected to its initial reference view $\widehat{V}(a)$ to calculate the current depth ratio ${\mathcal{R}}_{\widehat{V}(a)}(G_{a})$, which is compared with its initial depth ratio through VDRC to validate its own change ${\mathcal{P}}_{\widehat{V}(a)}(G_{a})=|\widehat{{\mathcal{R}}}_{\widehat{V}(a)}(G_{a})-{\mathcal{R}}_{\widehat{V}(a)}(G_{a})|/\widehat{{\mathcal{R}}}_{\widehat{V}(a)}(G_{a})$. Those splats not meeting the VDRC requirement will be excluded from the following splitting process. If $G_{a}$ is reliable by self-validation, the splitting is carried out, and VDRC continues to validate its two children $G_{b}$ and $G_{c}$ on their parent’s reference view $V(a)$ by Eq. 5. The valid child splat $G_{c}$ will regard its own initial reference view $\widehat{V}(c)$ inherited from $V(a)$ and initial depth ratio ^$\widehat{{\mathcal{R}}}_{\widehat{V}(c)}(G_{c})={\mathcal{R}}_{V(a)}(G_{c})$, which will be further used to verify whether $G_{c}$ is to be divided into new Gaussians in the next time of splitting.
@@ VI-C Normal Guided Splitting
$&\mu_{c}=(\widehat{\mu}_{c}-\mu^{\bot})\nabla I_{V(a)}(\bar{{\mathbf{x}}}_{c})+\mu^{\bot}\\ &\mu^{\bot}=\widehat{\mu}_{c}-((\widehat{\mu}_{c}-\mu_{a})\cdot N_{V(a)}(\bar{{\mathbf{x}}}_{a}))N_{V(a)}(\bar{{\mathbf{x}}}_{a}),\end{array}$ (6)
@@ VII Experiments
Table I: Quantitative comparison of our method to original 3DGS [1], GeoGaussian [15], Pixel-GS [17] and Mini-Splatting-D [14] on eight indoor scenes and seven outdoor scenes, with the number of Gaussian splats denoted by $\lvert\mathcal{G}\rvert$ in millions, and evaluation of rendering quality metrics with SSIM, PSNR in dB and LPIPS.
{tblr}
@@
cell12 = c=4, cell16 = c=4, cell110 = c=4, cell114 = c=4, cell118 = c=4, vline2,6,10,14,18 = -, hline1,3,11,18 = -, hline2 = 2-22, Methods & Original 3DGS GeoGaussian Pixel-GS Mini-Splatting-D Ours

### v2

@@ Abstract
publicationid: pubid: Copyright © 2024 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
@@ I Introduction
To better tackle both densification and positioning problems mentioned above, this paper presents a novel 3D Gaussian splatting framework for high-quality photorealistic novel view rendering, which we named GeoTexDensifier. Our framework innovatively proposes a densification strategy that enforces the reconstructed Gaussians to more accurately comply with the actual geometric structure of the scene. Moreover, our densified Gaussian model contains a sufficient number of splats to support fully textured regions, while maintaining sparse Gaussian splat distribution in weakly textured areas. In these ways, our novel 3DGS framework is able to produce high-quality 3D Gaussian models, as shown in the comparative example of “Church” in Fig. 1. The original 3DGS [1] has “over-reconstruction” problem that causes missing details in textured regions such as the carpet, while state-of-the-art (SOTA) works like GeoGaussian [15] improves the spatial structure of splat distribution and Pixel-GS [17] further densifies the splats in some textured regions like the carpet, but both of them still lack sufficient splat sampling to support texture details. Mini-Splatting [14] has sufficiently sampled splats but introduces over-densification situation in textureless regions, which might produce extra noisy Gaussians due to underconstrained visual ambiguity. In comparison, our GexTexDensifier delivers more accurate spatial distribution of Gaussian splat with less noise and more photorealistic novel view rendering effects according to the quantitative evaluation of the rendering results in the ground truth (GT) testing views on metrics of Structural SIMilarity (SSIM), Peak Signal to Noise Ratio (PSNR) in dB, and Learned Perceptual Image Patch Similarity (LPIPS). Our GeoTexDensifier framework carries out a geometry-aware splitting strategy to guide the splitting position sampling to fit the actual geometric surfaces of the scene with the help of normal prior, and filter out noisy splats with improper initial positions according to our Validation of Depth Ratio Change checking on relative depths provided by monocular depth prior. Meanwhile, a texture-aware densification strategy is adopted as auxiliary supplement for splitting of additional large Gaussians to produce a denser splat distribution in highly textured regions while remaining sparse in weakly textured areas, thereby preserving the overall high quality of Gaussian point cloud. The texture-aware densification and geometry-aware splitting strategies are fully integrated in the iterative optimization to get the final high-quality photorealistic 3DGS model. Experiments on Mip-NeRF 360 [6], Tanks and Temples [16] datasets and self-captured scenes verify the effectiveness and robustness of our GeoTexDensifier pipeline.
@@ II-C Gaussian Optimization with Geometry Prior
Original 3DGS expects Gaussian splats expected to grow along the real scene surfaces, which is however not always conformed to. Consequently, some approaches have attempted to optimize Gaussians by incorporating geometric priors like depths and normals. Similar to MonoSDF [33], DN-Splatter [34] employs mono-depth and mono-normal estimates to constrain the rendering loss. DNGaussian [35] introduces Hard and Soft Depth Regularization to optimize the positions and opacities of Gaussians, and employs Global-Local Depth Normalization to mitigate the sensitivity of scale-invariant loss to small depth errors. PGSR [36] optimizes Gaussians by rendering unbiased depths combined with single-view and multi-view geometric consistency losses. Besides applying depth and normal constraints, both GaussianRoom [37] and GSDF [38] utilize a learnable neural SDF field to guide the growth of splats, so as to simultaneously optimize both SDF and Gaussians. These geometry prior based methods rely on the geometric accuracy of the prior guidance, and are easily influenced by the incorrect depths or normals, especially when the truth depth scale is unavailable, while our approach only uses relative depth ratios for verifying newly generated splats, so as to relax the requirement on depth or scale accuracy.
@@ III System Overview
Suppose each scene has a set of multi-view RGB images captured with digital cameras as input, denoted by $\mathcal{I}=\{I_{t}|t=1,\dots,M\}$, where $M$ is the number of input images. Our GeoTexDensifier system is applied to the multi-view images as training views to robustly reconstruct an high-quality 3DGS model of the captured scene, which we defined as $\mathcal{G}$. Fig. 2 outlines the proposed high-quality 3DGS framework. In the initialization stage, COLMAP [39] is carried out first to estimate camera poses of all the input views denoted as $\{\mathbf{M}_{t}|t=1,\dots,M\}$, with sparse Structure-from-Motion (SfM) feature map points used to initialize a set of Gaussian splats. Meanwhile, a depth map is estimated for each input RGB image by ZoeDepth [40] with its normal map computed from depths, denoted as $\mathcal{D}=\{D_{t}|t=1,\dots,M\}$ and $\mathcal{N}=\{N_{t}|t=1,\dots,M\}$ respectively. After initialization finishes, the adaptive optimization stage follows to control the splat distribution by densification and pruning, while iteratively refining the Gaussian parameters to enforce their rasterized images to be consistent with the texture details observed in the training views. We use a texture-aware densification strategy to provide sufficient Gaussian splats in fully textured areas while maintaining the sparsity of Gaussians in regions with weak textures to ensure more accurate spatial distribution of Gaussian points. Additionally, the estimated depths are used as priors to guide the splitting of Gaussian splats to better conform to the actural tangential directions of the geometric surfaces in weakly textured regions and avoid the generation of isolated or scattered Gaussians caused by the ambiguity due to lack of visual features. The texture-aware densification approach together with the geometry-guided splitting strategy constitute our geometry-texture-aware densification framework for reconstructing high-quality photorealistic 3DGS model, which will be described in detail in the following sections.
@@ IV Preliminary
[…]
where ${\mathrm{C}}(\bar{{\mathbf{x}}})$ represents the color rendered at image pixel $\bar{{\mathbf{x}}}$, and $w_{k}(\bar{{\mathbf{x}}})$ is the rendering weight for $\alpha$-blending. $\bar{G}_{k}\sim\mathcal{N}(\bar{\mu}_{k},\bar{\Sigma}_{k})$ is the projected 2D Gaussian distribution of $G_{k}$ through a local affine approximation of perspective transformation $\mathbf{W}$, defined as $\bar{\mu}_{k}={\mathbf{W}}\mu_{k}$ and $\bar{\Sigma}_{k}=[{\mathbf{J}}{\mathbf{W}}\Sigma_{k}{\mathbf{W}}^{\top}{\mathbf{J}}^{\top}]_{1:2,1:2}$, with $\mathbf{J}$ the Jacobian form of $\mathbf{W}$ and $\cdot]_{1:2,1:2}$ taking the first two rows and columns as sub-matrix. Finally, by leveraging the differentiable rasterizer and comparing the rendered images to the training views, all attributes of the 3D Gaussians can be learned and optimized by minimizing the ${\mathcal{L}}_{1}$ loss combined with a D-SSIM loss.
@@
[…]
@@ V Texture-Aware Densification
An ideal densification strategy is to maintain dense splat distribution in strongly textured areas, while still leaving Gaussian splats relatively sparse in more weakly textured regions such as surfaces with pure colors. The blur split strategy introduced by Mini-Splatting [14] is adopted in our ADC process, which is although helpful for densification, but might cause the following side effects which do not actually conform to our ideal purpose: firstly, weakly textured regions often have Gaussians with large contribution areas, leading to over-splitting of splats to consequently generate too many inadequately optimized floaters which will influence the rendering results, and the second is that an excessive number of Gaussians in uniform splat distribution over the whole scene will cost too heavy memory consumption and large output file storage.
@@
which effectively reflects the contribution to whether a Gaussian point should be split by a hyperbolic tangent function normalized to $[0,1]$, which is activated by the image gradient whose gray values are normalized from $[0,255]$ to $[0,1]$, as shown in Fig. 3(a). $\alpha_{s}$ and $\beta_{s}$ are the linear regression coefficiencies which we set to 2020 and $0.16$ respectively for our experiments. Fig. 3(c) shows a computed weight map by Eq. 3. With the newly defined weight, we can redefine the maximal contribution area of each Gaussian $G_{i}$ to $I_{t}$ as $\hat{S}_{i}^{t}$, by modifying Eq. 2 to:
[…]
@@
Table I: The statistics of time in minutes (m), GPU peak memory (G. P. M.) consumption (GB), and rendering frame rate of the reconstructed 3DGS model in Frames Per Second (FPS) on our pipeline and Mini-Splatting-D [14] for the case “Bonsai” from Mip-NeRF 360 [6].
Methods Mini-Splatting-D Ours
Case Time G. P. M. FPS Time G. P. M. FPS
Bonsai 28m16s 6.61GB 132 42m31s 3.75GB 177
This texture-aware densification serves as an auxiliary strategy that supplements the original splitting strategy for a better ADC. Table I provides space/time resource statistics of training time, GPU peak memory (G. P. M.), and rendering frame rate (FPS) on our pipeline and Mini-Splatting-D [14], from which we can see that although our VDRC checking and normal guidance strategies which are described in section VI increase our training time complexity, our texture-aware splitting successfully suppresses the GPU memory usage, resulting in a smaller number of Gaussians to be rendered in a higher FPS.
@@ VI-A Reference View Selection and Updating
Figure 8: Two failure cases of “Horse” and “Truck” from Tanks and Temples dataset [16]. Each case contains a representative GT testing view and the rendered 3D Gaussian splats by our GeoTexDensifier.
@@ VI-B VDRC for Parent Splats
Before each new time of splitting, a certain number of iterations might shift some parent splats to improper locations. For example, $G_{a}$ might move from the door towards the wall, becoming an unreliable Gaussian like $G_{b}$, and will continue to generate a series of improper splats which will all meet the condition of VDRC. To better solve this issue, we apply VDRC to the parent Gaussians before each time we split them. Since the initial Gaussians are inherited from the SfM map points, we filter out the map points whose track lengths are less than $3$ or reprojection errors exceed $1$ pixel, to ensure the correctness of the initial splat locations. We also choose the image view with the smallest reprojection error denoted by $\widehat{V}(a)$ as the initial reference view of each Gaussian $G_{a}$, and add two additional attributes to record $\widehat{V}(a)$ and its initial depth ratio $\widehat{{\mathcal{R}}}_{\widehat{V}(a)}(G_{a})$. Before each time of splitting, each optimized parent Gaussian splat $G_{a}$ involved will be projected to its initial reference view $\widehat{V}(a)$ to calculate the current depth ratio ${\mathcal{R}}_{\widehat{V}(a)}(G_{a})$, which is compared with its initial depth ratio through VDRC to validate its own change ${\mathcal{P}}_{\widehat{V}(a)}(G_{a})=|\widehat{{\mathcal{R}}}_{\widehat{V}(a)}(G_{a})-{\mathcal{R}}_{\widehat{V}(a)}(G_{a})|/\widehat{{\mathcal{R}}}_{\widehat{V}(a)}(G_{a})$. Those splats not meeting the VDRC requirement will be excluded from the following splitting process. If $G_{a}$ is reliable by self-validation, the splitting is carried out, and VDRC continues to validate its two children $G_{b}$ and $G_{c}$ on their parent’s reference view $V(a)$ by Eq. 5. The valid child splat $G_{c}$ will regard its own initial reference view $\widehat{V}(c)$ inherited from $V(a)$ and initial depth ratio $\widehat{{\mathcal{R}}}_{\widehat{V}(c)}(G_{c})={\mathcal{R}}_{V(a)}(G_{c})$, which will be further used to verify whether $G_{c}$ is to be divided into new Gaussians in the next time of splitting.
@@ VI-C Normal Guided Splitting
[…]
@@ VII Experiments
Table II: Quantitative comparison of our method to original 3DGS [1], original 3DGS with depth prior [1], GeoGaussian [15], Pixel-GS [17] and Mini-Splatting-D [14] on Mip-NeRF 360 and Tanks and Temples datasets and two self-captured outdoor scenes, with the number of Gaussian splats denoted by $\lvert\mathcal{G}\rvert$ in millions, and evaluation of rendering quality metrics with SSIM, PSNR in dB and LPIPS.
@@
Methods Orig. 3DGS Orig. 3DGS + Depth Prior GeoGaussian Pixel-GS Mini-Splatting-D Ours
Cases $\lvert\mathcal{G}\rvert$ SSIM PSNR LPIPS $\lvert\mathcal{G}\rvert$ SSIM PSNR LPIPS $\lvert\mathcal{G}\rvert$ SSIM PSNR LPIPS $\lvert\mathcal{G}\rvert$ SSIM PSNR LPIPS $\lvert\mathcal{G}\rvert$ SSIM PSNR LPIPS $\lvert\mathcal{G}\rvert$ SSIM PSNR LPIPS
Bicyle (Mip-NeRF 360) 6.1 0.765 25.205 0.21 5.24 0.763 25.216 0.215 2.8 0.75 24.891 0.237 9.06 0.778 25.265 0.181 6.02 0.798 25.542 0.158 6.03 0.782 25.388 0.177
Garden 5.88 0.866 27.34 0.107 4.12 0.864 27.39 0.109 2.91 0.858 27.117 0.121 8.78 0.871 27.492 0.098 5.82 0.877 27.517 0.091 5.36 0.87 27.561 0.104
Stump 4.92 0.773 26.621 0.215 4.42 0.77 26.593 0.217 1.93 0.753 25.502 0.247 6.66 0.786 26.842 0.187 5.37 0.804 27.114 0.169 4.93 0.786 26.874 0.195
Treehill 3.77 0.633 22.523 0.325 3.37 0.634 22.525 0.325 1.91 0.629 22.058 0.344 7.97 0.634 22.21 0.276 4.86 0.641 22.222 0.261 3.73 0.635 22.379 0.293
Flowers 3.63 0.606 21.578 0.336 3.22 0.606 21.582 0.338 2.09 0.578 21.065 0.371 7.48 0.635 21.546 0.262 4.87 0.643 21.528 0.254 6.18 0.621 21.251 0.26
Room 1.54 0.919 31.414 0.219 1.37 0.919 30.982 0.218 0.62 0.901 29.908 0.251 2.57 0.922 31.579 0.208 3.93 0.928 31.576 0.187 2.55 0.927 31.93 0.195
Counter 1.21 0.909 28.993 0.2 1.13 0.907 28.874 0.2 0.56 0.889 27.899 0.235 2.57 0.915 29.186 0.183 3.78 0.913 28.728 0.171 1.74 0.916 29.216 0.179
Bonsai 1.26 0.942 31.991 0.203 1.15 0.94 32.219 0.204 0.76 0.925 31.026 0.225 2.08 0.947 32.554 0.191 3.76 0.947 31.957 0.174 1.83 0.954 32.952 0.161
Kitchen 1.82 0.926 31.036 0.127 1.6 0.924 30.999 0.128 0.73 0.908 29.207 0.148 3.19 0.931 31.702 0.119 3.67 0.933 31.802 0.114 2.09 0.931 31.698 0.12
Courtroom (Tanks & Temples) 2.99 0.818 23.621 0.217 3.23 0.806 23.788 0.218 1.47 0.797 23.152 0.265 5.18 0.818 23.719 0.215 5.24 0.792 21.73 0.219 4.78 0.829 24.287 0.199
Auditorium 0.69 0.881 24.083 0.247 0.77 0.871 24.793 0.264 0.25 0.867 23.765 0.292 1.21 0.883 24.351 0.244 4.8 0.839 21.506 0.274 1.64 0.892 25.24 0.234
Meetingroom 1.33 0.88 25.664 0.216 1.39 0.865 25.699 0.227 0.57 0.861 24.752 0.264 3.16 0.881 25.857 0.21 4.21 0.869 24.859 0.21 2.2 0.888 26.355 0.201
Church 2.24 0.829 22.946 0.247 2.61 0.804 22.553 0.266 1 0.811 22.355 0.279 3.61 0.833 23.317 0.243 4.86 0.818 21.89 0.241 4.08 0.845 23.678 0.217
Ballroom 3.42 0.807 23.902 0.188 3.33 0.813 24.236 0.183 1.23 0.788 23.31 0.141 6.07 0.808 23.843 0.181 4.61 0.806 23.49 0.179 4.45 0.815 23.99 0.17
Museum 5.34 0.774 21.22 0.235 4.49 0.788 21.956 0.222 2.42 0.76 20.43 0.176 9.62 0.769 21.15 0.239 6.09 0.77 20.624 0.231 4.76 0.778 21.527 0.226
Barn 1.03 0.817 26.828 0.259 1.13 0.815 26.921 0.256 0.39 0.788 25.54 0.24 4.22 0.84 27.234 0.215 3.88 0.862 27.455 0.175 4.02 0.856 27.472 0.18
Caterpillar 1.22 0.733 22.738 0.3 1.52 0.74 22.608 0.292 0.42 0.692 21.45 0.297 3.96 0.773 23.051 0.243 3.9 0.773 22.667 0.232 3.07 0.775 23.185 0.228
Courthouse 0.62 0.781 21.919 0.288 0.89 0.79 22.305 0.271 0.2 0.755 20.57 0.306 1.36 0.787 22.126 0.279 5.75 0.736 19.073 0.326 1.23 0.796 22.194 0.254
Family 2.31 0.844 24.024 0.191 2.23 0.848 24.289 0.182 1.12 0.807 22.460 0.164 3.86 0.847 23.942 0.174 4.21 0.856 24.27 0.154 2.07 0.848 24.218 0.187
Francis 0.81 0.881 27.188 0.265 0.87 0.883 27.798 0.267 0.36 0.86 25.13 0.162 2.02 0.889 27.521 0.249 3.89 0.888 26.203 0.238 1.27 0.888 27.375 0.245
Horse 1.32 0.866 23.481 0.187 1.49 0.874 24.249 0.175 0.72 0.85 22.73 0.139 2.39 0.869 23.744 0.167 3.91 0.87 23.849 0.152 1.18 0.862 23.613 0.187
Ignatius 3.45 0.732 21.245 0.264 2.96 0.736 21.817 0.259 1.56 0.704 20.460 0.234 5.92 0.744 21.59 0.235 4.75 0.747 21.273 0.217 4.46 0.744 21.781 0.228
Lighthouse 0.85 0.826 21.966 0.234 0.97 0.83 22.612 0.224 0.26 0.803 21.08 0.192 2.81 0.825 21.495 0.225 4.76 0.79 18.679 0.267 0.88 0.83 22.434 0.218
M60 1.75 0.887 26.839 0.195 1.58 0.889 27.361 0.193 0.75 0.87 26.040 0.151 5.04 0.896 27.329 0.164 4.14 0.895 26.961 0.153 2.42 0.895 27.17 0.163
Palace 0.66 0.741 19.936 0.359 0.84 0.745 20.398 0.342 0.23 0.725 19.51 0.358 1.90 0.746 20.014 0.346 7.28 0.666 15.467 0.444 0.85 0.751 20.352 0.326
Panther 1.96 0.894 27.734 0.185 1.73 0.894 27.753 0.185 0.96 0.882 27.04 0.122 5.14 0.899 27.705 0.158 4.41 0.871 25.571 0.184 1.79 0.896 27.675 0.177
Playground 2.32 0.801 24.209 0.267 2.03 0.805 24.684 0.262 1.15 0.775 23.3 0.224 4.17 0.832 24.829 0.217 4.56 0.871 25.571 0.184 3.21 0.835 24.853 0.197
Temple 0.76 0.788 20.021 0.292 1.03 0.804 21.072 0.27 0.29 0.772 19.630 0.27 3.38 0.789 19.993 0.272 5.46 0.7 15.755 0.358 0.91 0.795 20.483 0.266
Train 0.93 0.762 21.269 0.282 1.28 0.771 21.438 0.265 0.32 0.728 20.05 0.27 3.5 0.778 21.076 0.25 4.05 0.766 19.621 0.25 2.59 0.785 21.264 0.224
Truck 2.88 0.816 23.157 0.236 2.64 0.819 23.5 0.227 1.2 0.791 22.18 0.155 6.23 0.823 23.314 0.204 4.64 0.83 23.263 0.165 2.74 0.819 23.199 0.226
Wufeng Tower 4.69 0.775 23.073 0.292 5.21 0.781 23.159 0.286 1.62 0.729 21.351 0.355 11.98 0.777 22.848 0.267 5.86 0.768 22.495 0.273 10.34 0.8 23.402 0.232
Office Park 2.09 0.684 22.96 0.42 2.37 0.819 23.394 0.396 0.53 0.643 22.588 0.477 1.82 0.668 22.991 0.448 7.16 0.527 15.413 0.557 7.02 0.754 24.309 0.327
Table III: Ablation comparisons of only geometry-aware splitting (Geo.), only texture-aware densification (Tex.), texture-aware densification with VDRC (depth prior) (Tex. + VDRC), texture-aware densification with normal guidance (Tex. + Normal), and GeoTexDensifier (Tex. + Geo.) on Mip-NeRF 360 [6], with the number of Gaussian splats denoted by $\lvert\mathcal{G}\rvert$ in millions, and evaluation of rendering quality metrics with SSIM, PSNR in dB and LPIPS.
Methods Only Geo. (VDRC + Normal) Only Tex. Tex. + VDRC Tex. + Normal GeoTexDensifier (Tex. + Geo.)